OpenStudy (anonymous):

Naomi uses a credit card with a 25.9% APR, compounded monthly, to pay for a hotel stay totaling $971.20. She can pay $400 per month on the card. What will the total cost of this purchase be? $1,009.36 $1,222.74 $1,180.56 $971.20

OpenStudy (anonymous):

25.9/12 = 2.16% Compounding monthly interest

OpenStudy (anonymous):

Initial Amount: 971.20 Payments: 400 each month 971.20 * .0216 = 20.98 992.18 - 400 = 592.18 592.18 * .0216 = 12.79 604.97 - 400 = 204.97 204.97 * .0216 = Interest paid: 20.98 12.79 4.43 Interest combined: 38.2 Total Payment: 971.20 + 38.2 = $1009.40 Very close to the first answer, would this just be from my calculator and only bring over the first 3 numbers beyond the decimal point?

OpenStudy (fallenangelorchid):

ok answer is 1009.36, if you apply int to 1st month as well thats what you get

OpenStudy (anonymous):

Yeah I applied the interest to the initial amount owed aswell

OpenStudy (fallenangelorchid):

So in that case, your correct.

OpenStudy (anonymous):

Sweet, thanks for the second opinion :)

OpenStudy (fallenangelorchid):

Anytime lad.